[figure: 9.7 cm]

THE ANISOTROPY OF THE ELECTRICAL RESISTIVITY AND

THERMAL EXPANSION OF SINGLE CRYSTAL DyNi_{5}

E. Kačmarčiková^{a}, M. Reiffers^{b}, M. Mihalik^{b}, J. Kováč^{b}, A. Menovsky^{c},

J. Garcia Soldevilla^{d} and J.C. Goméz Sal^{d} aFaculty of Sciences, University of P.J.Šafárik, Park Angelinum 9, SK-041 54 Košice, Slovakia

bInstitute of Experimental Physics, SAS, Watsonova 47, SK-043 53 Košice, Slovakia

cDepartamento de Fisica, Universidad de Oviedo, E-33007 Oviedo, Spain dVan der Waals-Zeeman Institute, University of Amsterdam, 1018 XE Amsterdam, The Netherlands

The intermetallic compound DyNi_{5} crystallizes in hexagonal structure and undergoes the phase transition into antiferromagnetically ordered state. We report the results on the measurements of temperature dependence of the electrical resistivity ρ(T), thermal expansion L(T) and susceptibility χ(T) for single crystal sample in the temperature range 4.2 - 300 K. Residual resistivity ratios RRR = \frac{ρ(300K)}{ρ_{0}} along the main crystalographical axis vary from 3.5 to 4.7. We have observed the maximum in ρ(T) at T_{N} = 12.5 K which corresponds to the transition to the antiferromagnetically ordered state. The transition temperature is in an agreement with previous measurements. Taking into account Bloch-Grüneissen formula for electron-phonon contribution ρ_{ph}(T) to ρ(T) the anisotropic magnetic contribution ρ_{mag}(T) has been determined. We have observed the anisotropic behaviour of L(T) at low temperatures and the magnetic contributions to total L(T) have also been determined.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : M. Reiffers

Address for correspondence : Institute of Experimental Physics Watsonova 47 SK-043 53 Košice Slovakia

Email address : reiffers@kosice.upjs.sk